# Rectifying Conformity Scores for Better Conditional Coverage

### v1

@@ A.5 Comparison with simple baselines
These graphs provide some important insights:
@@
  • RCP in combination with any of these conformity score either maintains or improves conditional coverage and volume.
@@ B.3 Proof of Theorem 3
We will now proceed with the proof of Theorem 3, which verifies the marginal validity of our proposed approach. First, recall that $V\colon\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ is a conformity score function to which we apply a measurable transformation $(t,x)\in\mathbb{T}\times\mathcal{X}\to f_{t}^{-1}(x)$. For $x\in\mathcal{X}$, recall that RCP constructs the following prediction sets
@@ B.4 Proof of equality (10)
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{P}(\tilde{f}_{\varphi}^{-1}\circ V(X,Y)\leq\tau_{\star}(X)|X=x)\stackrel{{\scriptstyle(c)}}{{\geq}}1-\alpha,$ (31)
@@ B.5 Proof of Theorem 4
This section is dedicated to the proof of the conditional guarantee provided in Section 6. Along this section, we denote $\tilde{V}(x,y)=f_{\hat{\tau}(x)}^{-1}(V(x,y))$ and for any $t\in\mathbb{R}$, we denote
@@
$\epsilon_{\tau}(x)=\mathbb{P}\left({\tilde{f}_{\varphi}^{-1}(V(x,Y))\leq\hat{\tau}(x)\,|\,X=x}\right)-1+\alpha.$
@@
$1-\alpha+\epsilon_{\tau}(X_{n+1})-\alpha\mathrm{L}\times[{F_{\tilde{V}}(\varphi)}]^{n+1}\leq\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\,|\,X_{n+1}}\right)\\ \leq 1-\alpha+\epsilon_{\tau}(X_{n+1})+\mathrm{L}({\textstyle{1-\alpha+(n+1)^{-1}}})\times[{1-F_{\tilde{V}}(\varphi)}]^{n+1}.$
@@
$\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\,|\,X_{n+1}=x}\right)$ $=\int\mathbb{P}\left({f_{\tau(x)}^{-1}(V(x,Y))\leq t\,|\,X=x}\right)\textup{P}_{Q}(\mathrm{d}t)$
@@
00 0 $\leq F_{\tilde{V}\mid X=x}(\varphi)-F_{\tilde{V}\mid X=x}\circ F_{\tilde{V}}^{-1}\circ F_{\tilde{V}}(\varphi)$
@@
Since $U_{(k)}$ follows a beta distribution with parameters $(k,n+1-k)$, it follows that
@@ B.6 Pointwise control of $\epsilon_{\tau}$
Since $\mathcal{L}_{x}(t)$ is convex with Lipshitz continuous gradient, applying ([)Theorem 2.1.5]nesterov1998introductory, it follows that
@@ B.7.1 Step 1: intermediate results for Theorem 12
The following statement controls $\mathbb{P}({\sup_{v\in\mathbb{R}}\{{G(v)}\}\geq\epsilon})$. Its proof is similar to the extension of the Dvoretzky–Kiefer–Wolfowitz inequality provided in ([)Appendix B]plassier2024efficient.
@@ B.8 Asymptotic conditional validity
[…]
@@
Since $m$ is of the same order as $n$, which means that 0<lim $0<\liminf m/n\leq\limsup m/n<\infty$, it holds
@@
$01_{X\notin B_{\infty,h_{X}}}\mathds{1}_{X\in G}\xrightarrow[h_{X}\to 0]{}$ 0 .
@@ Appendix C Details on experimental setup
This section aims to provide addtional details on our experimental setup and implementation of the RCP algorithm.
@@
RCPMLP.  This variation reserves a part ($50\%$) of the original calibration set to train a quantile regression model for the $(1-\alpha)$ level quantile of the scores $V$. We again use a three hidden layers with 100 units per layers for that task. The remaining half of the calibration set forms the “proper calibration set” and is used to compute the conformal correction.
RCPlocal.  The local quantile regression variant is similar to the previous one, so we use the same splitting of the available calibration data. Since only one bandwidth needs to be tuned, we use a simple grid search on a log-scale grid in the interval $[10^{-3},1]$.

### v2

@@ A.7 Case of ellipsoidal prediction sets
The graphs on Figure 11 provide some important insights:
@@
  • RCP in combination with a better predictive model either maintains or improves conditional coverage and volume.
A.8 Improved data efficiency using cross-validation
As explained in Section 7, RCP requires to divide the calibration dataset $\mathcal{D}$ into two parts, one to estimate $\widehat{\tau}$, and one for SCP.
In this section, we consider a more data-efficient approach using the training dataset $\mathcal{D}_{\text{train}}$. Using $K$-fold cross-validation on $\mathcal{D}_{\text{train}}$, for each fold index $k$, we train a model on the K1K-1 remaining folds and evaluate the conformity score on the fold $k$. This yields a dataset $\mathcal{D}_{\tau}$ of size $|\mathcal{D}_{\text{train}}|$ with inputs and their associated conformity scores based on which $\widehat{\tau}$ is estimated. This also removes the need to split the calibration dataset. An additional model is fitted on the complete training data set $\mathcal{D}_{\text{train}}$ to produce the non-rectified conformity scores.
Figure 12 shows a comparison of learning $\widehat{\tau}$ on half the calibration dataset (cal), or using 10-fold cross-validation (CV). The cross-validation approach yields improved worst-slab coverage on most datasets. This improved conditional coverage comes at the computational cost of training $K$ additional models.
Figure 12: Worst-slab coverage of RCP with $\widehat{\tau}$ trained on half the calibration dataset (cal) or using 10-fold cross-validation (CV).
A.9 Comparison with CPCG
[…]
Table 4: Comparison of worst-slab coverage on multi-output datasets.
$_{\text{cal}}-PCP RCP$$_{\text{CV}}-$PCP CPCG-PCP $_{\text{cal}}-DCP RCP$$_{\text{CV}}-$DCP CPCG-DCP
households 0.825 0.905 0.899 0.888 0.853 0.891 0.900 0.900
scm20d 0.830 0.892 0.891 0.897 0.877 0.877 0.868 0.899
rf1 0.731 0.830 0.877 0.838 0.715 0.863 0.827 0.872
scm1d 0.758 0.882 0.895 0.910 0.756 0.902 0.896 0.882
meps_21 0.739 0.874 0.904 0.881 0.789 0.881 0.879 0.905
meps_19 0.762 0.875 0.867 0.880 0.788 0.884 0.889 0.878
meps_20 0.731 0.842 0.871 0.890 0.719 0.880 0.884 0.892
house 0.835 0.895 0.903 / 0.817 0.878 0.906 /
bio 0.784 0.860 0.900 0.887 0.774 0.879 0.880 0.880
blog_data 0.770 0.877 0.893 0.886 0.749 0.844 0.888 0.888
calcofi 0.810 0.889 0.888 / 0.828 0.885 0.892 /
taxi 0.837 0.885 0.884 0.881 0.846 0.872 0.879 0.879
Table 5: Comparison of computational time (in seconds) on multi-output datasets.
$_{\text{cal}}-PCP RCP$$_{\text{CV}}-$PCP CPCG-PCP $_{\text{cal}}-DCP RCP$$_{\text{CV}}-$DCP CPCG-DCP
households 0.258 0.604 104 8840 0.00759 0.531 104 8164
scm20d 2.63 4.22 772 6409 0.0182 0.852 766 6012
rf1 0.667 1.35 340 10682 0.00836 0.348 339 9674
scm1d 2.27 3.57 1209 4971 0.0133 0.867 1205 4692
meps_21 0.236 0.607 581 6283 0.0123 0.261 581 6031
meps_19 0.272 0.515 493 6411 0.0123 0.184 492 6128
meps_20 0.255 0.520 621 7147 0.0119 0.238 621 7032
house 0.315 0.594 1034 / 0.0159 0.327 1033 /
bio 0.630 1.22 3161 79422 0.0279 0.782 3163 63178
blog_data 0.752 0.850 1119 41192 0.0336 0.155 1121 43289
calcofi 0.699 1.02 456 / 0.0356 0.200 455 /
taxi 0.680 1.17 866 77828 0.0269 0.276 866 70139
@@ B.3 Proof of Theorem 1
We will now proceed with the proof of Theorem 1, which verifies the marginal validity of our proposed approach. First, recall that $V\colon\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ is a conformity score function to which we apply a measurable transformation $(t,x)\in\mathbb{T}\times\mathcal{X}\to f_{t}^{-1}(x)$. Recall that RCP constructs the following prediction sets for $x\in\mathcal{X}$
@@ B.4 Proof of equality (9)
$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{P}(\tilde{f}_{\varphi}^{-1}\circ V(X,Y)\leq\tau_{\star}(X)|X=x)\stackrel{{\scriptstyle(c)}}{{\geq}}1-\alpha,$ (30)
@@ B.5 Proof of Theorem 2
This section is devoted to the proof of the conditional guarantee given in Section 6. In this section, we denote $\tilde{V}(x,y)=f_{\hat{\tau}(x)}^{-1}(V(x,y))$ and for each $t\in\mathbb{R}$, we denote
@@
$\epsilon_{\tau}(x)=\mathbb{P}\left({\tilde{f}_{\varphi}^{-1}(V(x,Y))\leq\widehat{\tau}(x)\,|\,X=x}\right)-1+\alpha.$
@@
$1-\alpha+\epsilon_{\tau}(x)-\alpha\mathrm{L}\times[{F_{\tilde{V}}(\varphi)}]^{n+1}\leq\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\,|\,X_{n+1}=x}\right)\\ \leq 1-\alpha+\epsilon_{\tau}(x)+\mathrm{L}({\textstyle{1-\alpha+(n+1)^{-1}}})\times[{1-F_{\tilde{V}}(\varphi)}]^{n+1}.$
@@
$\mathbb{P}\left({Y\in\mathcal{C}_{\alpha}(X)\,|\,X=x}\right)$ $=\int\mathbb{P}\left({f_{\tau(x)}^{-1}(V(x,Y))\leq t\,|\,X=x}\right)\textup{P}_{Q}(\mathrm{d}t)$
@@
$0$ $\leq F_{\tilde{V}\mid X=x}(\varphi)-F_{\tilde{V}\mid X=x}\circ F_{\tilde{V}}^{-1}\circ F_{\tilde{V}}(\varphi)$
@@
Since $U_{(k)}$ follows a beta distribution with parameters (k,n+1k)(k,n+1-k), it follows that
@@ B.6 Pointwise control of $\epsilon_{\tau}$
Since $\mathcal{L}_{x}(t)$ is convex with Lipshitz continuous gradient, applying Theorem 2.1.5 from (Nesterov, 1998), it follows that
@@ B.7.1 Step 1: intermediate results for Theorem 7
The following statement controls $\mathbb{P}({\sup_{v\in\mathbb{R}}\{{G(v)}\}\geq\epsilon})$. Its proof is similar to the extension of the Dvoretzky–Kiefer–Wolfowitz inequality provided in Appendix B of (Plassier et al., 2024).
@@ B.8 Asymptotic conditional validity
$A_{n}=\left\{{x\in\mathcal{X}\colon F_{\tilde{V}\mid X=x}\circ F_{\tilde{V}}^{-1}\text{ is $4^{-1}\rho[{F_{\tilde{V}}(\varphi)}]^{-n}\wedge[{1-F_{\tilde{V}}(\varphi)}]^{n+1}$-Lipschitz}}\right\}.$
@@
Since $m$ is of the same order as $n$, which means that $0<\liminf m/n\leq\limsup m/n<\infty$, it holds
@@
$1_{X\notin B_{\infty,h_{X}}}\mathds{1}_{X\in G}\xrightarrow[h_{X}\to 0]{}0.$
@@ Appendix C Details on Experimental Setup
This section aims to provide additional details on our experimental setup and implementation of the RCP algorithm.
@@
RCPMLP.  This variation reserves a part ($50\%$) of the original calibration set to train a quantile regression model for the $(1-\alpha)$-level quantile of the scores $V$. We again use a three hidden layers with 100 units per layers for that task. The remaining half of the calibration set forms the “proper calibration set” and is used to compute the conformal correction.
RCPlocal.  The local quantile regression variant is similar to the previous one, so we use the same splitting of the available calibration data. Since only one bandwidth needs to be tuned, we use a simple grid search on a log-scale grid in the interval $10^{-3},1]$.